# Protein Structure Tokenization via Geometric Byte Pair Encoding

### v1

@@ Appendix I Hyperparameter Settings
Appendix K Algorithmic Details
@@
5:  Run RMSD_Partition (Alg. 6) to obtain medoids and $c:\mathcal{O}(\kappa^{\star})\!\to\!\{1,\dots,K_{\kappa^{\star}}\}$.
6:  Define $\mathcal{A}_{\kappa^{\star}}=\{\Pi^{(\kappa^{\star})}_{j}\}_{j=1}^{K_{\kappa^{\star}}}$ (medoid spans’ internal-parameter tuples).
@@
13:(Optional periodic global glue opt). If mode=all and the schedule triggers, apply GlueOptAll (Alg. 10) to all modified backbones; recompute keys for their adjacent pairs, and for every affected key $k$, perform the same erase/insert priority update with $\rho(k)=\mathbf{1}[k\notin\mathrm{dom}(\mathcal{V})]$. If $\textsc{Front}(\mathcal{D})$ then exposes a recurring key $(\rho=0)$ promoted by glue refinement, immediately re-invoke Step (no new clustering).
@@
0:GeoBPE vocab $\mathcal{V}=\{\kappa\mapsto\mathcal{A}_{\kappa}\}$ with key introduction order $(\kappa^{(1)},\dots,\kappa^{(S)})$; medoids $\mathcal{A}_{\kappa}=\{\Pi^{(\kappa)}_{j}\}_{j=1}^{K_{\kappa}}$; glue quantizers $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$ with bin centers $\{\mu_{b}^{\theta}\}_{b=1}^{B_{\theta}}$, $\{\mu_{b}^{\omega}\}_{b=1}^{B_{\omega}}$, $\{\mu_{b}^{\phi}\}_{b=1}^{B_{\phi}}$

### v2

@@ Appendix I Performance Across Protein Fold Types
Appendix M Algorithmic Details
Additional notation for algorithms.
We reuse all geometric and GeoBPE notation from Secs. 3.13.2. For convenience we collect the additional symbols that appear only inside the algorithmic pseudocode.
$\mathcal{S},\mathcal{A}$
Set of motif (or motif–pair) occurrences and its sampled subset used by RMSD_Partition (Alg. 6); each $u\in\mathcal{S}$ indexes a motif $\mathcal{M}^{(t_{u})}_{i_{u}{:}k_{u}}$.
$\mathcal{S}_{3},\mathcal{S}_{2}$
Collections of interior and terminal bond–residue occurrences used to build residue-level codebooks $\mathcal{A}_{3},\mathcal{A}_{2}$ (Algo. 18).
$\widehat{\mathcal{M}}$, $c(\cdot)$
Medoid set and assignment map returned by RMSD_Partition, used as inputs to glue-optimization routines (Algos. 19,12).
$\mathcal{D}^{(\star)},\ \mathcal{O}^{(\star)}(\kappa)$
Single-backbone geo-pair map and occurrence sets for a new backbone $t^{\star}$ during tokenization (Algo. 10).
$\Sigma,\Sigma_{\mathrm{med}}$
Token dictionary used for geometric language modeling (Algos. 16, 17); $\Sigma_{\mathrm{med}}$ contains only motif tokens; $\Sigma$ also includes glue angle tokens.
$\mathrm{id}_{\mathrm{med}},\mathrm{id}_{\mathrm{bin}}$
Integer maps assigning token IDs to motif medoids $(\kappa,j)$ and to glue-angle bins $(\mathrm{type},b)$, respectively.
$x^{(\tau)}$
Token sequence encoding backbone $t_{\tau}$ obtained by alternating motif and glue-bin tokens (Algo. 17).
$h_{i}^{\uparrow},c_{i}^{\uparrow},\bar{h}_{i},\bar{c}_{i}$
Upward and downward TreeLSTM states at node $i$ in the up–down encoder (Algos. 1315).
$z^{\mathrm{res}}_{\tau,i},\ z^{\mathrm{prot}}_{\tau}$
Final residue-level and protein-level embeddings produced by the up–down encoder on the merge hierarchy $\mathcal{F}^{(\tau)}$ (Algo. 15).
@@
5:  Run RMSD_Partition (Alg. 6) to obtain medoids and $c:\mathcal{O}(\kappa^{\star})\!\to\!\{1,\dots,K_{|\kappa^{\star}|}\}$.
6:  Define $\mathcal{A}_{\kappa^{\star}}=\{\Pi^{(\kappa^{\star})}_{j}\}_{j=1}^{K_{|\kappa^{\star}|}}$ (medoid spans’ internal-parameter tuples).
@@
13:(Optional periodic global glue opt). If mode=all and the schedule triggers, apply GlueOptAll (Alg. 12) to all modified backbones; recompute keys for their adjacent pairs, and for every affected key $k$, perform the same erase/insert priority update with $\rho(k)=\mathbf{1}[k\notin\mathrm{dom}(\mathcal{V})]$. If $\textsc{Front}(\mathcal{D})$ then exposes a recurring key $(\rho=0)$ promoted by glue refinement, immediately re-invoke Step (no new clustering).
@@
Algorithm 10 Tokenize — use learned GeoBPE vocabulary to tokenize new backbone
0: New backbone $t^{\star}$ (length NN); learned residue codebooks $\mathcal{A}_{3},\mathcal{A}_{2}$; learned vocabulary $\mathcal{V}$ whose pair-keys are ordered by training insertion, written
$\mathrm{Order}(\mathcal{V})=\big\langle\kappa_{1},\kappa_{2},\ldots,\kappa_{|\mathcal{V}|}\big\rangle,$
with each $\kappa_{\ell}$ mapped to its fixed prototype set $\mathcal{A}_{\kappa_{\ell}}=\{\Pi^{(\kappa_{\ell})}_{j}\}_{j=1}^{K_{|\kappa_{\ell}}|}$; quantizers $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$; optional glue mode $\in\{\textit{none},\textit{each},\textit{all}\}$ and (if all) a period $P$.
0: Tokenized segmentation $\mathcal{P}^{(\star)}$ and merge hierarchy $\mathcal{F}^{(\star)}$ for $t^{\star}$.
1:Per-residue init (no new clustering). Set $\mathcal{P}^{(\star)}\!\leftarrow\!(\mathcal{M}^{(t^{\star})}_{1{:}1},\ldots,\mathcal{M}^{(t^{\star})}_{N{:}N})$. Assign each residue motif to the nearest element of $\mathcal{A}_{3}$ (interior) or $\mathcal{A}_{2}$ (terminal) by Kabsch-aligned RMSD, and overwrite its internal parameters accordingly.
2:Initialize hierarchy. Let $\mathcal{F}^{(\star)}$ be a binary forest whose leaves (in order) are $\{\mathcal{M}^{(t^{\star})}_{i{:}i}\}_{i=1}^{N}$; its frontier equals $\mathcal{P}^{(\star)}$.
3:Optional one-time global glue. If mode==all, run GlueOptAll (Alg. 12) on $t^{\star}$ once; snap $(\theta^{C\!N\!CA},\omega,\phi)$ to $Q$.
4:Build single-structure geo-pair map. Compute $\mathcal{D}^{(\star)}\!\leftarrow\!\textsc{BinHelper}\big(t^{\star},\mathcal{P}^{(\star)},Q\big)$ (Alg. 11), which maps any geo-pair key $\kappa$ to its occurrence set $\mathcal{O}^{(\star)}(\kappa)$ on $t^{\star}$. (Uses ComputeGeoKey with raw medoid internals and quantized boundary glue.)
5:Apply learned merges in training order (no new keys).
6:for $s=1$ to $|\mathcal{V}|$} do
7:  $\kappa\leftarrow\kappa_{s}$  (the $s$-th key in $\mathrm{Order}(\mathcal{V})$).
8:  if $\kappa\notin\mathcal{D}^{(\star)}$ then
9:   continue
10:  end if
11:  Assign prototypes (no clustering). For each $(\mathcal{L},\mathcal{R})\in\mathcal{O}^{(\star)}(\kappa)$ with $\mathcal{L}=\mathcal{M}^{(t^{\star})}_{p{:}q}$, $\mathcal{R}=\mathcal{M}^{(t^{\star})}_{q+1{:}r}$, set
$c(\mathcal{L},\mathcal{R})\;=\;\arg\min_{j\in\{1,\ldots,K_{|\kappa|}\}}\mathrm{RMSD}\big(\mathcal{M}^{(t^{\star})}_{p{:}r},\ \Pi^{(\kappa)}_{j}\big)$.
[…]
  1. 1.
    Form $\widetilde{\mathcal{M}}=\mathcal{M}^{(t^{\star})}_{p{:}r}$ and overwrite its internals by $\Pi^{(\kappa)}_{j}$.
  2. 2.
    Update segmentation $\mathcal{P}^{(\star)}$: replace $(\mathcal{L},\mathcal{R})$ by $\widetilde{\mathcal{M}}$.
  3. 3.
    Update hierarchy $\mathcal{F}^{(\star)}$: add a parent for span [p:r][p{:}r] with left child the node of $\mathcal{L}$ and right child the node of $\mathcal{R}$; update the frontier so it equals the new $\mathcal{P}^{(\star)}$.
  4. 4.
    If mode==each and the boundary $p{-}1\!\to p$ exists, apply GlueOpt (Alg. 19); snap its three angles to $Q$.
13:  Maintain the single-structure map. For each merged $(\mathcal{L},\mathcal{R})\in S^{(\star)}$, update $\mathcal{D}^{(\star)}$ locally: remove the occurrence of $\kappa$; decrement keys of neighbors $(\mathcal{L}^{-},\mathcal{L})$ and $(\mathcal{R},\mathcal{R}^{+})$ (when defined); insert the new neighbor keys (,~)($\mathcal{L}^{-}$$\widetilde{\mathcal{M}}$) and (~,+)($\widetilde{\mathcal{M}}$$\mathcal{R}^{+}$) using ComputeGeoKey.
14:  Optional periodic global glue. If mode==all and $s\bmod P=0$, run GlueOptAll on $t^{\star}$; then recompute keys adjacent to changed boundaries via ComputeGeoKey and refresh their occurrences in $\mathcal{D}^{(\star)}$.
15:end for
16:return $\mathcal{P}^{(\star)}$ and $\mathcal{F}^{(\star)}$.
Algorithm 11 BinHelper — build geo-pair map for one backbone
0: Backbone $t^{\star}$ with current segmentation $\mathcal{P}^{(\star)}$; quantizers $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$.
[…]
1: Initialize $\mathcal{D}^{(\star)}\leftarrow\emptyset$.
2:for each adjacent pair $(\mathcal{L},\mathcal{R})$ in $\mathcal{P}^{(\star)}$ do
3:  $\kappa\leftarrow\textsc{ComputeGeoKey}(\mathcal{L},\mathcal{R})$ using:
  • Raw medoid internals for $\mathcal{L}$ and $\mathcal{R}$ (as assigned in initialization or prior merges);
  • boundary glue $(\theta^{C\!N\!CA},\omega,\phi)$ snapped by $Q$.
4:  Insert the occurrence $(\mathcal{L},\mathcal{R})$ into $\mathcal{O}^{(\star)}(\kappa)$.
5:end for
6:return $\mathcal{D}^{(\star)}$.
@@
0:GeoBPE vocab $\mathcal{V}=\{\kappa\mapsto\mathcal{A}_{\kappa}\}$ with key introduction order $(\kappa^{(1)},\dots,\kappa^{(S)})$; medoids $\mathcal{A}_{\kappa}=\{\Pi^{(\kappa)}_{j}\}_{j=1}^{K_{|\kappa|}}$; glue quantizers $Q_{\theta^{C\!N\!CA}},Q_{\omega},Q_{\phi}$ with bin centers $\{\mu_{b}^{\theta}\}_{b=1}^{B_{\theta}}$, $\{\mu_{b}^{\omega}\}_{b=1}^{B_{\omega}}$, $\{\mu_{b}^{\phi}\}_{b=1}^{B_{\phi}}$